A Bibliography of Typography Publications

Karl Berry 135 Center Hill Rd. Plymouth, MA 02360 USA

E-mail: karl@cs.umb.edu (Internet)

12 August 2014 Version 1.39

Title word cross-reference

Aided [Car86, KW79]. Alphabets [Sol83, Tsc85]. Anatomy [Law90]. Art [Sol86]. Aspects [WBTW90]. Atlas [SB88]. Automatic [Joh87, Sch88].

Based [Sch88].

Century [Car87]. Classic [Sol83]. Communication [CDM85]. Complete [Sol83]. Computer [Car86, KW79]. Computer-Aided [Car86, KW79]. Conventional [Joh87]. Credit [Tra86]. Cubics [PS83]. Curve [PS83, Sch88]. Curve-fitting [PS83]. Curves [Sch88].

Dear [Rey82]. Design [Ada86, Bri90, Car86, CDM85, Gou77, Har75, KW79, Rug87, Sch88]. Designers [Car87]. Designing [Cra71]. Designs [Joh87]. Dialogue [Ada86]. Digital [Ada86]. Digitial [Joh87].

Environment [Rug87]. Extensible [Rug87].

Face [Byr91]. Fitting [Sch88, PS83]. Fly [URW91a]. Font [Ada86, And92]. Fonts [Sol83]. Form [CDM85, Har75]. Format [Div92]. Forms [Ada86].

Hand [Sch88]. Hand-Sketched [Sch88]. Hudson [McL80]. hz [URW91b]. hz-program [URW91b].

Intellifont [Div92]. Interactive [Rug87, Sch88]. Inversions [Kim89].

Just [Dür17].

Kerning [URW91a]. Kq [URW91c].

Kq-program [URW91c].

Letter [KW, Kin76, KW79]. Lettering [Har75, Tsc85]. Letters [Ada86, Dür17, Har75, Tra86].

Making [Gou77]. Manual [McL80].

Metrics [And92]. Minutes [Bri90]. Model [Joh87]. My [Rey82].

Optical [Kin76].

Paragon [Rug87]. Parametric [PS83]. Phoenix [Sch88]. Phototypesetting [URW91c, URW91b]. Piecewise [PS83]. PRINT [Byr91]. Printing [Kin76]. program [URW91c, URW91b].

Roman [Sol83]. Runemeister [Rey82].

Scalable [Div92]. Scaling [Joh87]. Shaping [Dür17]. Sketched [Sch88]. Skill [Har75]. Space [Bri90]. Spacing [KW, Kin76]. Studies [Gou77]. System [Sch88]. Systems [Kin76].

Technology [Joh87]. Ten [Bri90]. Text [WBTW90]. Thames [McL80]. Treasury [Tsc85]. Twentieth [Car87]. Type [Bri90, Car87, Cra71, Gou77, Joh87]. Typeface [Div92, Car86, Law90, Rug87]. Typeforms [SB88]. Typographic [CDM85]. Typography [McL80, Sol86]. Typologia [Gou77].

URW [URW91c, URW91b]. Use [Har75].

Visible [WBTW90].

References

Adams:DFL-86

[Ada86] Debra Anne Adams. A dialogue of forms: Letters and digital font design. Master’s thesis, Massachusetts Institute of Technology, MIT Microreproduction Lab, 77 Massachusetts Avenue, room 014-0551, Cambridge, MA 02139. $46.35, September 1986.

Andre:IRISA-92-676

[And92] Jacques André. Font metrics.

Technical Report 676, IRISA, Campus de Beaulieu, F-35042 Rennes cedex, September 1992.

Briem:HST-90

[Bri90] Hunnlauger SE Briem. How to space a type design in ten minutes. Lecture notes for the Type 90 conference, Oxford, August– September 1990.

Byrne:Print-91-9

[Byr91] Chuck Byrne. PRINT’s new face. Print, pages 45–53, September/ October 1991. ISSN 0032-8510.

Carter:CTD86

[Car86] Kathleen Anne Carter. Computer-Aided Typeface Design. PhD thesis, U. Cambridge, Corn Exchange St., Cambridge CB2 3QG, U.K., May 1986.

Carter:TCT87

[Car87] Sebastian Carter. Twentieth Century Type Designers. Taplinger Publishing Co., New York, 1987. ISBN 0-8008-7916-3. 168 pp. LCCN Z250.A2 C3717 1987.

Carter:TDF85

[CDM85] Rob Carter, Ben Day, and Philip B. Meggs. Typographic Design: Form and Communication. Van Nostrand Reinhold Company Inc., New York, 1985. ISBN 0-442-26166-7. LCCN 85-667.

Craig:DT71

[Cra71] James Craig. Designing with Type. Watson-Guptill Publications, New York, 1971. ISBN 0-8230-1320-0. LCCN 70-159564.

Agfa:ISF90

[Div92] Agfa Division. Intellifont Scalable Typeface Format. Miles Inc., 90 Industrial Way, Wilmington, MA 01887, 2nd edition, 1992.

Durer:JSL-17

[Dür17] Albrecht Dürer. Of the Just Shaping of Letters. Grolier Club, New York, 1917. LCCN N 6888 D8A125E, Z239.2 .R64 1917d, A23-671.

Goudy:T77

[Gou77] Frederic W. (Frederic William)

Goudy. Typologia: Studies in Type Design & Type Making. University of California Press, Berkeley, 2nd edition, 1977. ISBN 0-520-03278-0 (paperback). xviii + 170 pp. LCCN Z250 .G6 1977.

Harvey:LD75

[Har75] Michael Harvey. Lettering Design: Form & Skill in the Design and Use of Letters. Barre Publishers, Barre, MA, 1975. ISBN 0-517-52744-8. 159 pp. LCCN NK3600 .H32 1976.

Johnson:MAS87

[Joh87] Bridget Lynn Johnson. A model for automatic scaling of type designs for conventional and digitial technology. Master’s thesis, Rochester Institute of Technology, May 1987.

Kim:I89

[Kim89] Scott Kim. Inversions. W.H.

Freeman, San Francisco, New York, 1989.

Kindersley:OLS76

[Kin76] David Kindersley. Optical Letter Spacing for New Printing Systems. Wynkyn de Worde Society, distributed by Lund Humphries Publishers Ltd., 26 Litchfield St. London WC2, 1976.

Kindersley:LS-27266/77

[KW] David G. Kindersley and Neil E. Wiseman. Letter spacing. British Patent Application 27266/77.

Kindersley:PW-1979-oct-12

[KW79] David Kindersley and Neil Wiseman. Computer-aided letter design. Printing World, pages 12–17, October 1979. CODEN PROWAE. ISSN 0032-8715.

Lawson:AT90

[Law90] Alexander Lawson. Anatomy of a Typeface. David R. Godine, Boston, 1990.

McLean:THT80

[McL80] Ruari McLean. The Thames and Hudson Manual of Typography. Thames and Hudson, Inc., New York, 1980.

Plass:CG-17-229

[PS83] Michael Plass and Maureen Stone. Curve-fitting with piecewise parametric cubics. Computer Graphics, 17(3):229–239, July 1983. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Reynolds:MDR82

[Rey82] Lloyd Reynolds. My Dear Runemeister. Taplinger Publishing Co., Inc, New York, NY, USA, 1982.

Ruggles:PIE87

[Rug87] Lynn Elizabeth Ruggles. Paragon, An Interactive, Extensible, Environment for Typeface Design. PhD thesis, University of Massachusetts, Amherst, September 1987.

Sutton:AT88

[SB88] James Sutton and Alan Bartram.

An Atlas of Typeforms. Chartwell Books, Secaucus, second edition, 1988. ISBN 1-55521-340-5.

Schneider:PIC-88

[Sch88] Philip J. Schneider. Phoenix: An interactive curve design system based on the automatic fitting of hand-sketched curves. Master’s thesis, U. Washington, 1988.

Solo:CRA83

[Sol83] Dan X. Solo. Classic Roman Alphabets: 100 Complete Fonts. Dover pictorial archive series. Dover Publications, New York, 1983. ISBN 0-486-24517-9. 100 pp. LCCN Z250.5.R64 C57 1983.

Solomon:AT86

[Sol86] Martin Solomon. The Art of Typography. Watson-Guptill Publications, New York, 1986.

Tracy:LC86

[Tra86] Walter Tracy. Letters of Credit.

David R. Godine, Boston, 1986. US$27.50.

Tschichold:TAL88

[Tsc85] Jan Tschichold. A Treasury of Alphabets and Lettering. Omega Books Ltd., 1 West Street, Ware, Hertfordshire, 1985. ISBN 1-85007-007-5.

URW:kerning

[URW91a] URW. Kerning on the fly. In URW Non-Plus-Ultra Typography series, Harksheider Str. 107, D-2000 Hamburg 65, 1991.

URW:phototypesetting-hz

[URW91b] URW. Phototypesetting with the URW hz-program. In URW Non-Plus-Ultra Typography series, Harksheider Str. 107, D-2000 Hamburg 65, 1991.

URW:phototypesetting-Kq

[URW91c] URW. Phototypesetting with the URW kq-program. In URW Non-Plus-Ultra Typography series, Harksheider Str. 107, D-2000 Hamburg 65, 1991.

Watt:AVP-90-309

[WBTW90] R. J. Watt, J. Bock, H. Thimbleby, and A. Wilkins. Visible aspects of text. In Applying Visual Psychophysics to User Interface Design, pages 309–325. Lavenham, May 1990.